OpenStudy (fanduekisses):

Finding the critical numbers of f(theta)=2cos(theta) + sin^2(theta )

OpenStudy (fanduekisses):

\[f(\theta)=2\cos(\theta) + \sin^2(\theta)\]

OpenStudy (fanduekisses):

\[f'(\theta)=-2\sin(\theta) + 2\sin(\theta)*\cos(\theta) = 0\]

OpenStudy (fanduekisses):

\[f'(\theta)= -2\sin(\theta)(1-\cos(\theta))= 0\]

OpenStudy (fanduekisses):

\[\theta= 2\pi \]

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you have done all the work, solve \[\sin(\theta)=0\] and also \[1-\cos(\theta)=0\]

OpenStudy (fanduekisses):

theta = pi, 2pi